What are vinyl halides? Can you give the structures of some examples?

Grade:11

3 Answers

Mahima Kanawat
5 years ago
Dear student 
Vinyl halides are haides attached to vinyl group. 
Vinyl group is CH2=CH -
If halogen is chlorinec is vinyl chloride
Ex CH2=CH - Cl
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT 
kunta gouthamreddy
5 years ago
 
Vinyl halides are haides attached to vinyl group. 
Vinyl group is CH2=CH -
If halogen is chlorinec is vinyl chloride
Ex CH2=CH - Cl
kunta gouthamreddy
5 years ago
In organic chemistry, a vinyl halide is a compound with the formula CH2=CHX (X = halide). The term vinyl is often used to describe any alkenyl group. For this reason, alkenyl halides with the formula RCH=CHX are sometimes called vinyl halides. From the perspective of applications, the dominant member of this class of compounds is vinyl chloride, which is produced on the scale of millions of tons per year as a precursor to polyvinyl chloride.[1] Polyvinyl fluoride is another commercial product. Related compounds include vinylidene chloride and vinylidene fluoride.